Find the point of intersection of the pair of straight lines.

y= −8x − 9
−y= 9x + 7

Answer by mananth(16946)

8 x + 1 y = -9 .............1
-9 x -1 y = 7 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 1
8 x 1 y = -9
-9 x -1 y = 7
Add the two equations
-1 x = -2
/ -1
x = 2
plug value of x in (1)
8 x + 1 y = -9
16 + 1 y = -9
1 y = -9 -16
1 y = -25
y = -25

m.ananth@hotmail.ca